# Hyper-Transforming Latent Diffusion Models

### v2

@@ 4.5 Data completion
Figure 6: Conditional inpainting results on CelebA. The second and third rows present two different conditional samples generated by our LDMI for the missing regions at the centre of the ground truth. Our model successfully reconstructs coherent structures and fine details, demonstrating its ability to generate diverse and plausible completions.
Lastly, we demonstrate our model’s performance in data completion tasks. In our experiments, we evaluate the model’s performance on both structured and unstructured missing data patterns. We include random pixel masks for irregular regions of missing data and structures missing patterns of occlusion. For each test case, we show the original image, the masked input and our model’s imputed output.
The results in Figure 6, demonstrate that our approach can effectively leverage context from available input to generate plausible completions for missing areas. For facial images from CelebA, our model successfully reconstructs missing facial features while maintaining consistency with visible regions in the input. Similarly, for ImageNet samples, the model generates contextually appropriate textures and structures that seamlessly integrate with surrounding areas.
@@ 5 Conclusion
We introduced Latent Diffusion Models of Implicit Neural Representations (LDMI), a novel framework that combines the expressiveness of INRs with the generative power of Latent Diffusion Models (LDMs). Our proposed hyper-transforming approach enables efficient adaptation of pre-trained LDMs to INR-based generation without extensive retraining. Experimental results across CelebA, ImageNet, ShapeNet Chairs, and ERA5 demonstrate the effectiveness of our method, achieving competitive performance in both quantitative and qualitative evaluations.
Beyond empirical findings, this work establishes a novel paradigm for resolution-agnostic generative modelling by integrating INRs with latent diffusion, paving the way for future advances in generative modelling where resolution independence and structural consistency are crucial.
@@ A.1 Autoencoding-based models
When opting by $\beta$-VAE Higgins et al. [2017] for the first stage of LDMs Rombach et al. [2022], a weak regularization encourages accurate likelihoods $p_{\lambda}({\bm{y}}|{\bm{z}})$ parameterized by the decoder $D_{\lambda}(\cdot)$ while preserving latent expressiveness. This is achieved by assuming $p({\bm{z}})=\mathcal{N}(\bm{0},\bm{I})$ and maximizing the Evidence Lower Bound (ELBO):
@@
In contrast, VQ-VAE replaces the continuous latent representation with a discrete codebook. During training, the encoder output is quantized using the closest embedding from the codebook, enabling a more structured representation. The objective becomes:
@@ A.2.1 Inference
The reverse posterior density defined in Equation (21) is no longer Markovian and coincides with the inference model proposed in DDIM Song et al. [2021b], where it is demonstrated that faster sampling can be achieved without retraining, simply by redefining the generative process as:
The reverse posterior density defined in Equation (21) is no longer Markovian and coincides with the inference model proposed in DDIM Song et al. [2021b], where it is demonstrated that faster sampling can be achieved without retraining, simply by redefining the generative process as:
@@ Appendix B Training LDMI
     Compute likelihood loss: $\mathcal{L}_{\text{HT}}(\phi)$ using Eq. 15
@@
     Compute ELBO loss: $\mathcal{L}_{\text{VAE}}(\phi,\psi)$ using Eq. 11
@@
     Compute DDPM loss: $\mathcal{L}_{\text{DDPM}}$ using Eq. 14
@@ C.1 Scalability of LDMI
A key strength of our approach lies in its parameter efficiency and scalability when compared to alternative methods. As shown in Table 3, while previous approaches such as GASP [Dupont et al., 2022a] or VAMoH [Koyuncu et al., 2023] require substantial hypernetwork parameters (25.7M) to generate relatively small INR weights (50K), LDMI achieves superior performance with only 8.06M hypernetwork parameters while generating 330K INR weights for a 5-layer network.
@@ C.3 Reconstructions at multiple resolutions
Figures 8 and 9 present reconstructions from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. Reconstructions are rendered at resolutions ranging from ×$\times 0.125$ to ×$\times 4$, showcasing the model’s ability to maintain spatial coherence and fine details even under extreme upsampling. Notably, such flexible resolution control is not supported by baseline methods like GASP due to architectural limitations. Furthermore, datasets of this complexity are not even addressed by other reconstruction-capable baselines such as VAMoH [Koyuncu et al., 2023], Functa [Dupont et al., 2022a], or Spatial Functa [Bauer et al., 2023].
@@ Appendix D Experimental details
This section details the hyperparameter settings used in training LDMI across different datasets and configurations. We provide an overview of key model components, including the encoder, decoder, and diffusion model, as well as dataset-specific settings.
Table 5 lists the hyperparameters for all models used in our experiments. It covers both stages of our generative framework: (i) the first-stage autoencoder—either a VQ-VAE or VAE, depending on the dataset—and (ii) the second-stage latent diffusion model. The table includes architectural choices such as latent dimensionality, diffusion steps, attention resolutions, and optimization parameters (e.g., batch size, learning rate). It also details the HD decoder, tokenizer, Transformer modules, and INR architecture across different modalities. For image-based experiments, we additionally employ perceptual and adversarial losses to improve sample quality, following Esser et al. [2021]. All models were trained using NVIDIA H100 GPUs.
@@
first_stage_model VAE VAE VQ-VAE VQ-VAE VAE VAE

### v3

@@ 4.5 Data Completion
We evaluate our model on structured completion tasks, where parts of the input are masked with predefined patterns. For each case, we show the masked input and two samples drawn from the posterior. As shown in Figure 7, the model uses visible context to produce realistic, coherent outputs. On CelebA-HQ $256\times 256$, it reconstructs missing features—such as eyes, mouth, or hairline—while preserving consistency with the unmasked regions. These results highlight its ability to generalize from partial observations.
@@
Figure 7: Conditional inpainting results on CelebA-HQ at $256\times 256$. The second and third rows present two conditional samples generated by our LDMI for the missing regions at the centre.
@@ 5 Conclusion
We introduced Latent Diffusion Models of Implicit Neural Representations (LDMI), a framework that combines the expressiveness of INRs with the generative power of Latent Diffusion Models. Our hyper-transforming approach allows efficient adaptation of pre-trained LDMs to INR-based generation without retraining the full model. Experiments on CelebA-HQ, ImageNet, ShapeNet Chairs, and ERA5 demonstrate strong performance in both quantitative and qualitative evaluations.
Importantly, LDMI overcomes limitations of prior hypernetworks through a modular, probabilistic Transformer decoder that supports scalable, expressive INR generation. While MLP-based methods struggle with dense parameter outputs and Transformer-based ones are typically deterministic, our design enables probabilistic modelling and efficient synthesis. Integrating INRs with latent diffusion yields a flexible, resolution-agnostic framework—advancing generative tasks requiring structural consistency, fidelity, and scalability.
Acknowledgements
This research was supported by the Villum Foundation through the Synergy project number 50091, and by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (MLLS, grant NNF20OC0062606).
Ignacio Peis acknowledges support by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429). Batuhan Koyuncu is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Autoencoding-based Models
When opting by $\beta$-VAE (Higgins et al., 2017) for the first stage of LDMs (Rombach et al., 2022), a weak regularization encourages accurate likelihoods $p_{\lambda}({\bm{y}}|{\bm{z}})$ parameterized by the decoder $D_{\lambda}(\cdot)$ while preserving latent expressiveness. This is achieved by assuming $p({\bm{z}})=\mathcal{N}(\bm{0},\bm{I})$ and maximizing the Evidence Lower Bound (ELBO):
@@
In contrast, VQVAE replaces the continuous latent representation with a discrete codebook. During training, the encoder output is quantized using the closest embedding from the codebook, enabling a more structured representation. The objective becomes:
@@ A.2.1 Inference
The reverse posterior density defined in Equation (21) is no longer Markovian and coincides with the inference model proposed in DDIM (Song et al., 2021a), where it is demonstrated that faster sampling can be achieved without retraining, simply by redefining the generative process as:
@@ Appendix B Training LDMI
    Compute likelihood loss: $\mathcal{L}_{\text{HT}}(\phi)$ using Equation (15)
@@
    Compute ELBO loss: $\mathcal{L}_{\text{VAE}}(\phi,\psi)$ using Equation (11)
@@
    Compute DDPM loss: $\mathcal{L}_{\text{DDPM}}$ using Equation (14)
@@ C.1 Scalability of LDMI
A key strength of our approach lies in its parameter efficiency and scalability when compared to alternative methods. As shown in Table 3, while previous approaches such as GASP (Dupont et al., 2022b) or VAMoH (Koyuncu et al., 2023) require substantial hypernetwork parameters (25.7M) to generate relatively small INR weights (50K), LDMI achieves superior performance with only 8.06M hypernetwork parameters while generating 330K INR weights for a 5-layer network.
@@ C.3 Reconstructions at Multiple Resolutions
Figures 9 and 10 present reconstructions from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. Reconstructions are rendered at resolutions ranging from ×$\times 0.125$ to ×$\times 4$, showcasing the model’s ability to maintain spatial coherence and fine details even under extreme upsampling. Notably, such flexible reconstruction is not supported by baseline methods like GASP due to architectural limitations. Furthermore, datasets of this complexity are not even addressed by other reconstruction-capable baselines such as VAMoH (Koyuncu et al., 2023), Functa (Dupont et al., 2022b), or Spatial Functa (Bauer et al., 2023).
@@ Appendix D Experimental Details
This section outlines the hyperparameter settings used to train LDMI across datasets. We describe key components, including the encoder, decoder, diffusion model, and dataset settings. Table 5 summarizes the hyperparameters used in our experiments, covering both stages of the generative framework: (i) the first-stage autoencoder—either a VQVAE or $\beta$-VAE, depending on the dataset—and (ii) the second-stage latent diffusion model. It lists architectural choices such as latent dimensionality, diffusion steps, attention resolutions, and optimization settings (e.g., batch size, learning rate), along with details of the HD decoder, tokenizer, Transformer modules, and INR architecture. For image-based tasks, we use perceptual and adversarial losses to improve sample quality, following Esser et al. (2021). All models were trained on NVIDIA H100 GPUs.
@@
first_stage_model $\beta$-VAE $\beta$-VAE VQVAE VQVAE $\beta$-VAE $\beta$-VAE